OpenStudy (anonymous):

what is x? Log3(x+7) - Log3(x) = 2 and how did you get it?

OpenStudy (anonymous):

\[\log_{3} \frac{x+7}{x} = 2\] \[\frac{x+7}{x} = 3^{2}\]